Calculus I, Fall 2014 (http://www.math.nthu.edu.tw/~wangwc/)

Brief answer to selected problems in Homework 12

1. Section 6.1: problem 10: ∫_{−1}^{1} ^{\frac{(2\sqrt{1−y^{2}})^{2}}{2}} dy = \frac{8}{3}